OpenStudy (anonymous):

What is the next term in the sequence 2, 8, 32, 128, ...? 256 512 768 1024

OpenStudy (mathmale):

Hey, P-G, step back a moment and ask yourself: how do we get from 2 to 8? Then, how do we get from 8 to 32? We're looking for a pattern that we will summarize in a simple formula for the terms of this sequence.

OpenStudy (anonymous):

it would be 512. the pattern is the number gets multiplied by 4 to get the next number.

OpenStudy (zale101):

this seems like it's a geometric sequence because the sequence multiplies by 4 and geometric sequence use multiplication between the sequences. 2, *4 8, *4 32, *4 128, *4 ...?

OpenStudy (the_fizicx99):

This is clearly a geometric sequence as you can see that they're not increasing in a linear form, so because this is geometric you have to find the ratio, So you can divide any number by the previous number, so..

OpenStudy (anonymous):

512

OpenStudy (mathmale):

So, who can come up with a simple formula that works to predict the nth term of this sequence?

OpenStudy (the_fizicx99):

Mee!!

OpenStudy (mathmale):

Go riiiiight ahead! Thanks, Zale, for putting a name (geometric) to this sequence.

OpenStudy (anonymous):

http://www.wolframalpha.com

OpenStudy (anonymous):

they gave sequence when i plugged in numbers

OpenStudy (triciaal):

2^1, 2^3, 2^5, 2^7, 2^9

OpenStudy (the_fizicx99):

f(n) = f(n – 1) • r, where n > 0 The rate is 4 Suppose you want to find the 5th term (you do >.> ) f(5) = f(5-1) * 4 f(5) = f(4) * 4 f(5) = 128

OpenStudy (the_fizicx99):

512 * typo >.<

OpenStudy (the_fizicx99):

f(4) = 128 128 * 4 = 512 :>

OpenStudy (triciaal):

2^2n+1

OpenStudy (mathmale):

As Zale has correctly pointed out, this is a geometric sequence. Have any of you seen the formula a*r^n (or a*r^(n-1)? Could that be used to write a simple formula with which to predict the nth value of the sequence?

OpenStudy (the_fizicx99):

You can also use, f(n) = f(n – 1) • r Yes I've seen that formula :)

OpenStudy (mathmale):

that's a recursion formula; each term depends on the value of the preceding one. Can anyone come up with a formula that's NOT a recursion formula?

OpenStudy (the_fizicx99):

f(n) = f(1) • r^n-1

OpenStudy (mathmale):

Yes, or f(n)=2*(4)^(n-1). Agreed? With what value should n begin?

OpenStudy (the_fizicx99):

Your asking me? Or paris?